# Breaking the HBM Bit Cost Barrier: Domain-Specific ECC for AI Inference Infrastructure

### v1

@@ Abstract
High-Bandwidth Memory (HBM) delivers exceptional bandwidth and energy efficiency for AI workloads, but its high cost per bit, driven in part by stringent on-die reliability requirements, poses a growing barrier to scalable deployment. This work explores a system-level approach to cost reduction by eliminating on-die ECC and shifting all fault management to the memory controller. We introduce a domain-specific ECC framework combining large-codeword Reed–Solomon (RS) correction with lightweight fine-grained CRC detection, differential parity updates to mitigate write amplification, and tunable protection based on data importance. Our evaluation using LLM inference workloads shows that, even under raw HBM bit error rates up to $10^{-3}$, the system retains over 78% of throughput and 97% of model accuracy compared with systems equipped with ideal error-free HBM. By treating reliability as a tunable system parameter rather than a fixed hardware constraint, our design opens a new path toward low-cost, high-performance HBM deployment in AI infrastructure.
@@ I Introduction
High-Bandwidth Memory (HBM) has become a foundational element of modern AI computing systems, delivering the high memory bandwidth required to sustain the throughput of large-scale inference and training workloads. However, HBM remains significantly more expensive than conventional DRAM in terms of \$/GB, at least by a factor of 5 to 10, which increasingly threatens the scalability and economic viability of future AI infrastructure. Fundamentally, there are only two viable paths to address this cost challenge: (i) reducing the volume of data stored in HBM through compression, quantization, or algorithmic sparsification [1]; or (ii) reducing the manufacturing cost of HBM itself. While the first path has attracted extensive research, this work explores the second, approaching it from a system architecture perspective.
@@
The central idea of this work is to reduce HBM cost by substantially relaxing raw reliability requirements, shifting fault tolerance from HBM stack to the system level. This principle is well established in storage systems: hard disk drives (HDDs) and solid-state drives (SSDs) routinely tolerate elevated raw bit error rate (BER) by strengthening system-level fault tolerance mechanisms such as error-correcting codes (ECC). Unlike these general-purpose media, our focus is domain-specific: HBM deployed for AI inference. Two factors motivate this direction. First, AI inference is expected to dominate both compute and memory usage in future AI infrastructure, making it economically viable to pursue hardware and system-level customization. Second, inference workloads exhibit structural properties, such as sequential access patterns and unequal bit criticality, that expose a large, underexplored design space for HBM ECC solutions.
@@
Effectively preserving data integrity under relaxed HBM raw reliability requires significantly stronger error correction, and the only viable path is to increase ECC codeword length, from today’s typical 16B or 32B to much larger blocks such as 512B or 2KB. This is because ECC strength improves exponentially with codeword length. However, deploying large-codeword ECC introduces two critical challenges. First, small random memory accesses, which may still occur despite the predominantly sequential nature of AI inference, can suffer from severe read and write amplification. Second, large-size ECC requires high-complexity decoding hardware, significantly increasing the silicon area and power consumption of the memory controller. To address these limitations, we introduce two complementary techniques: (i) We propose a hybrid ECC architecture that pairs large-size Reed–Solomon (RS) codes with fine-grained Cyclic Redundancy Check (CRC), allowing the system to opportunistically bypass full RS decoding and thereby mitigate read/write amplification. (ii) We can selectively omit ECC protection for low-sensitivity data (e.g., mantissa fields) to reduce decoding complexity and energy without compromising inference quality. Together, these techniques enable practical adoption of large-codeword ECC in AI inference memory systems, unlocking cost reductions through reliability relaxation. Using large language model (LLM) inference as a case study, our evaluation shows that, even under very high raw HBM bit error rates (e.g., up to $10^{-3}$), the proposed approach can sustain throughput and inference accuracy comparable to systems equipped with ideal, error-free HBM. It is our hope that this work will inspire further research on system- and workload-aware reliability co-design, enabling scalable, cost-effective memory solutions for future AI infrastructure.
@@ II Background and Motivation
HBM achieves exceptional bandwidth through two key architectural features: a 3D stack of DRAM dies interconnected via dense through-silicon vias (TSVs), and a wide, high-speed interface that connects directly to the processor die within the same package. This tight integration enables thousands of parallel I/O wires, resulting in aggregate bandwidths of multiple terabytes per second. However, this same integration imposes significant manufacturing complexity and yield sensitivity, driving the cost per bit to at least 5$\sim$10$\times$ higher than conventional DRAM. As AI workloads become increasingly memory-bound, this cost disparity emerges as a critical obstacle to building scalable, cost-efficient AI infrastructure.
A major contributor to HBM cost is the stringent reliability requirement imposed during die testing and binning. To meet these requirements, each DRAM die integrates short-length on-die ECC applied to each 16B or 32B word. In addition to on-die ECC, HBM interfaces also permit the host to attach a 2B CRC to each 32B word. This CRC is generated and verified by the host-side memory controller and transmitted alongside the data for end-to-end integrity checking. While this combination of on-die ECC and host-managed CRC provides modest fault coverage, it suffers from several fundamental limitations. First, the use of short codewords limits the correction strength and scalability of the protection, making it inadequate for handling the reliability variability introduced by advanced process scaling. Second, both the ECC logic and redundancy budget are hardwired into the DRAM silicon at tape-out, eliminating the possibility of adapting to workload-specific resilience or tuning protection strength dynamically.
To increase yield and reduce manufacturing cost, a more flexible, system-driven reliability model is needed. One promising approach is to shift fault-tolerance responsibility from the HBM die to the host-side memory controller, thereby relaxing raw reliability requirements at the HBM stack level. This architectural shift enables the practical deployment of large-codeword ECC, which offers significantly stronger protection at the same redundancy overhead. For example, as shown in Fig. 1, increasing the codeword size from 32B to 512B or 2KB can improve the tolerable raw bit error rate (BER) by over fivw orders of magnitude, making it possible to qualify lower-cost, higher-yield DRAM dies without compromising end-to-end data integrity.
@@
Implementing large-codeword ECC directly on DRAM die is impractical due to its lack of high-performance logic circuit resource. A natural alternative is to migrate all fault management to the off-die, host-side memory controller that can support the implementation of large-size ECC encoder/decoder and flexible ECC configurations. The sequential access patterns of LLM inference workloads naturally amortize the cost of large-codeword ECC, keeping steady-state bandwidth overhead low.
@@ III Proposed Solutions
This work centers around two simple ideas: (1) applying large-size ECC to substantially relax HBM raw reliability constraints, thereby enabling lower-cost DRAM manufacturing; and (2) leveraging the inherent error resilience of AI inference workloads to reduce the implementation overhead of such ECC. Achieving this requires shifting ECC responsibilities from HBM stack to host-side memory controller, as illustrated in Fig. 2.
Figure 2: Overview of the proposed design framework.
@@ III-A Hybrid ECC Design
In current HBM product, on-die ECC is implemented using short RS codes, chosen for their robustness to both random and burst errors, and applied to 16B or 32B words. This work preserves the use of RS codes but extends them to operate over much larger codeword sizes (e.g., 512B or 2KB) under host control. Meanwhile, we also reuse another key feature of the existing HBM interface: the ability for host processors to append a 2B CRC to each 32B data chunk. Instead of relying on CRC solely for basic integrity checking, we integrate it into a hybrid RS/CRC scheme, where per-chunk CRCs serve as low-latency filters that detect errors and avoid triggering full RS decoding. This opportunistic use of CRC dramatically reduces read and write amplification, without modifying the HBM physical interface or access granularity.
To organize large-size RS codeword efficiently, we exploit the inherent parallelism of HBM stacks (e.g., each HBM3E stack includes 16 independent channels, each 128 bits wide). We stripe each RS codeword across $s$ such channels to maximize bandwidth and parallel access, as illustrated in Fig. 3. Specifically, each RS codeword contains a number of 32B user data chunks and one or multiple 32B parity chunks. Each 32B chunk, whether data or parity, is appended with its own 2B CRC, forming a 34B unit. These CRC-augmented chunks are striped sequentially across the $s$ channels. Because each chunk is self-contained with its own CRC, this structure supports fine-grained error detection with minimal latency and bandwidth overhead. The design aligns naturally with HBM’s native access granularity and requires no modification to the physical interface.
Figure 3: Placement of CRC-enhanced RS codewords over multiple HBM channels.
@@
Random Write: Fig. 5 illustrates the procedure for handling random write requests. For random writes that update $k$ 32B chunks within a single RS codeword, the memory controller performs a CRC-guided opportunistic partial update to avoid full read-modify-write operations. First, it fetches the current version of the $k$ target chunks along with all the $r$ RS parity chunks. If all CRCs on the fetched chunks pass, the controller proceeds with a partial update as follows. Assume each RS codeword protects a total of $m$ 32B chunks. We construct two $m\times$32B vectors ${\bf D}_{old}$ and ${\bf D}_{new}$: one containing the old values of the $k$ chunks in their original positions and zeros elsewhere, and another containing the new values of those same $k$ chunks (with zeros in the remaining positions). These two sparse vectors are independently RS-encoded to generate two sets of $r$ parity chunks $\text{RS}({\bf D}_{old})$ and $\text{RS}({\bf D}_{new})$. Because RS code is a linear block code, the difference between them captures the net parity update due to the modified data chunks. Accordingly, the new RS parity ${\bf D}_{new}$ is computed by XORing the original parity ${\bf P}_{old}$ with the difference of the two RS-encoded results:
$\mathbf{P}_{\text{new}}=\mathbf{P}_{\text{old}}\oplus\text{RS}(\mathbf{D}_{\text{new}})\oplus\text{RS}(\mathbf{D}_{\text{old}}).$
@@
Sequential Read/Write: Sequential read and write requests are handled in a much simplifier manner. For sequential reads, we directly fetch the entire RS codeword, including data and parity, and perform RS decoding without first checking CRCs. At high BER, errors are likely, and skipping CRC avoids redundant checks. The RS decoder can naturally early-terminate if no errors are present, minimizing latency and energy. For sequential writes, the controller performs RS encoding and per-chunk CRC encoding in a single pass, then writes the full hybrid RS/CRC codeword directly to HBM.
@@ III-B Importance‑Adaptive ECC Protection
As discussed above, implementing large-codeword RS decoders incurs substantial silicon cost, particularly under multi-TB/s HBM throughput constraints. To mitigate this burden, we propose importance-adaptive ECC protection, which selectively bypasses error protection for less critical data. Many AI inference workloads, especially those using floating-point formats like BF16 or FP8, exhibit inherent numerical resilience: mantissa bits can tolerate occasional bit flips with minimal impact on inference accuracy. By applying ECC only to the most sensitive portions of the data, we significantly reduce RS decoder workload, ECC-related traffic, and energy consumption.
To support this functionality efficiently, we adopt a bit-plane–oriented in-memory data placement scheme. Consider a block of $m$ numerical values $\{x_{1},\dots,x_{m}\}$, each $n$ bits wide. Each value $x_{j}$ is represented as a bit vector $[b_{j,n-1},b_{j,n-2},\dots,b_{j,0}]$, where $b_{j,i}\in\{0,1\}$ is the $i$-th bit of $x_{j}$. The $i$-th bit-plane is defined as $P_{i}=\{b_{1,i},b_{2,i},\dots,b_{m,i}\}$, i.e., all bits at position $i$ across the block. Let $\mathcal{S}\subseteq\{0,\dots,n-1\}$ denote the set of critical planes that require protection, and let $\gamma=\frac{|\mathcal{S}|}{n}$ denote the protected-plane ratio, with $0\leq\gamma\leq 1$. Only bits in planes $\mathcal{S}$ are processed through CRC and RS encoding; all other bits bypass the ECC logic entirely. Accordingly, we can reduce the RS decoder silicon cost by approximately $1-\gamma$.
@@ IV-A Impact of Codeword Size on Inference Throughput
These trends are driven by two antagonistic effects: (i) Amplification of random access. A CRC fault escalates random small chunks request to an entire codeword fetch or a read-modify-write, whose traffic penalty grows with the codeword length. (ii) Parity amortization. Sequential requests always transfer the entire codeword if the CRC fails. Enlarging the codeword length spreads the parity bytes over more useful data (i.e., higher code rate), which will improve the effective bandwidth.
@@ IV-B Impact of Random Access Intensity
We next evaluate the sensitivity of inference throughput to the fraction of random memory accesses under a high raw BER of $10^{-3}$. Fig. 7 shows token throughput as the random-access ratio increases from 0% to 10% for six RS codeword lengths. Under purely sequential traffic (0% random), throughput increases consistently with codeword size, rising from 13.90 tokens/s (64B) to 18.05 tokens/s (2048B). This trend highlights the benefit of longer codewords in amortizing ECC overhead: larger codewords have higher code rates, reducing redundancy and improving effective memory bandwidth.
However, introducing even a small fraction of random access disrupts this advantage. At 2% random access, throughput drops across all cases, with larger codewords most affected. The 64B case decreases slightly to 13.85 tokens/s (–0.4%), while the 2048B case drops more sharply to 14.26 tokens/s (–21.0%). At 10% random access, 64B still maintains 13.64 tokens/s (–1.9%), while 256B and 2048B fall to 11.87 (–19.2%) and 7.31 tokens/s (–59.5%), respectively. This pattern reflects the impact of access amplification. When a CRC failure occurs on a 34B chunk, it escalates into a full-codeword read or read-modify-write, and the associated overhead scales with codeword size. Consequently, in workloads with even modest random activity, moderate codeword lengths (256B–512B) offer the best trade-off between ECC efficiency and access amplification, while very large codewords become increasingly inefficient.

### v2

@@ Abstract
High-Bandwidth Memory (HBM) delivers exceptional bandwidth and energy efficiency for AI workloads, but its high cost per bit, driven in part by stringent on-die reliability requirements, poses a growing barrier to scalable deployment. This work explores a system-level approach to cost reduction by eliminating on-die ECC and shifting all fault management to the memory controller. We introduce a domain-specific ECC framework combining large-codeword Reed–Solomon (RS) correction with lightweight fine-grained CRC detection, differential parity updates to mitigate write amplification, and tunable protection based on data importance. Our evaluation using LLM inference workloads shows that, even under raw HBM bit error rates up to $10^{-3}$, the system retains 78% of throughput while maintaining at least 97% PIQA accuracy and 94% MMLU accuracy relative to error-free HBM. By treating reliability as a tunable system parameter rather than a fixed hardware constraint, our design opens a new path toward low-cost, high-performance HBM deployment in AI infrastructure.
@@ I Introduction
High-Bandwidth Memory (HBM) has become a foundational element of modern AI computing systems, delivering the high memory bandwidth required to sustain the throughput of large-scale inference and training workloads. However, HBM remains significantly more expensive than conventional DRAM in terms of \$/GB, at least by a factor of 5 to 10 [1], which increasingly threatens the scalability and economic viability of future AI infrastructure. Fundamentally, there are only two viable paths to address this cost challenge: (i) reducing the volume of data stored in HBM through compression, quantization, or algorithmic sparsification [2]; or (ii) reducing the manufacturing cost of HBM itself. While the first path has attracted extensive research, this work explores the second, approaching it from a system architecture perspective.
@@
The central idea of this work is to reduce HBM cost by substantially relaxing raw reliability requirements, shifting fault tolerance from HBM stacks to the system level. This principle is well established in storage systems: hard disk drives (HDDs) and solid-state drives (SSDs) routinely tolerate elevated raw bit error rate (BER) by strengthening system-level fault tolerance mechanisms such as error-correcting codes (ECC) [3]. Unlike these general-purpose media, our focus is domain-specific: HBM deployed for AI inference. Two factors motivate this direction. First, AI inference is expected to dominate both compute and memory usage in future AI infrastructure, making it economically viable to pursue hardware and system-level customization. Second, inference workloads exhibit structural properties, such as sequential access patterns and unequal bit criticality, that expose a large, underexplored design space for HBM ECC solutions.
@@
Effectively preserving data integrity under relaxed HBM raw reliability requires significantly stronger error correction, and the only viable path is to increase ECC codeword length, from today’s typical 16B or 32B  [4] to much larger blocks such as 512B or 2KB. This is because ECC strength improves exponentially with codeword length. However, deploying large-codeword ECC introduces two critical challenges. First, small random memory accesses, which may still occur despite the predominantly sequential nature of AI inference, can suffer from severe read and write amplification. Second, large-size ECC requires high-complexity decoding hardware, significantly increasing the silicon area and power consumption of the memory controller. To address these limitations, we introduce two complementary techniques: (i) We propose a hybrid ECC architecture inspired by concatenated codes [5] that pairs large-size Reed–Solomon (RS) codes with fine-grained Cyclic Redundancy Check (CRC), allowing the system to opportunistically bypass full RS decoding and thereby mitigate read/write amplification. (ii) We can selectively omit ECC protection for low-sensitivity data (e.g., mantissa fields) motivated by unequal error protection [6] to reduce decoding complexity and energy without compromising inference quality. Together, these techniques enable practical adoption of large-codeword ECC in AI inference memory systems, unlocking cost reductions through reliability relaxation. Using large language model (LLM) inference as a case study, our evaluation shows that, even under very high raw HBM bit error rates (e.g., up to $10^{-3}$), the proposed approach can sustain throughput and inference accuracy comparable to systems equipped with ideal, error-free HBM. It is our hope that this work will inspire further research on system- and workload-aware reliability co-design, enabling scalable, cost-effective memory solutions for future AI infrastructure.
@@ II Background and Motivation
[…]
A major contributor to HBM cost is the stringent reliability requirement imposed during die testing and binning. In HBM3, each DRAM die integrates short-length on-die ECC at 16B/32B granularity, and the interface permits the host to attach a 2B CRC to each 32B word [4]. The CRC is generated and verified by the host-side controller and transmitted with the data for end-to-end checking. While this combination improves field reliability, it has inherent limits: short codewords cap correction strength and scalability, and both the ECC logic and redundancy are fixed in the DRAM silicon, preventing workload-aware tuning as process scaling increases raw error variability.
To increase yield and reduce manufacturing cost, a more flexible, system-driven reliability model is needed. The core lever is to use larger ECC codewords at the same redundancy, which sharply improves correction strength. As a motivating example, Fig.1 models RS codes at a fixed code rate (16/17): starting from 32B (on-die ECC granularity in current HBM [4]), increasing the codeword size to 2KB raises the tolerable raw bit error rate (BER) by more than five orders of magnitude. This illustrates the headroom available from codeword scaling alone, independent of where ECC is implemented.
@@
Implementing such large-codeword ECC directly on DRAM dies is impractical due to limited high-performance logic resources and tight area/power budgets. A practical alternative is to move fault management to the host-side memory controller, which can host high-throughput RS encoders/decoders and support flexible ECC configurations. Critically, the largely sequential access patterns of LLM inference amortize large-codeword overheads in steady state, making controller-based ECC an effective path to relax raw HBM reliability while preserving end-to-end integrity.
@@ III Proposed Solutions
This work centers around two simple ideas: (1) applying large-size ECC to substantially relax HBM raw reliability constraints, thereby enabling lower-cost DRAM manufacturing; and (2) leveraging the inherent error resilience of AI inference workloads to reduce the implementation overhead of such ECC. Achieving this requires shifting ECC responsibilities from HBM stacks to host-side memory controller, as illustrated in Fig. 2.
Figure 2: Comparison of HBM ECC Architectures. (a) The state-of-the-art system uses small, independent on-die ECC codewords within each HBM channel. (b) Our proposed system shifts all fault management to the host controller’s ECC logic, and use large, cross-channel RS codewords on low-cost HBM stacks.
@@ III-A Hybrid ECC Design
In current HBM products, on-die ECC is implemented using short RS codes, chosen for their robustness to both random and burst errors, and applied to 16B or 32B words. This work preserves the use of RS codes but extends them to operate over much larger codeword sizes (e.g., 512B or 2KB) under host control. Meanwhile, we also reuse another key feature of the existing HBM interface: the ability for host processors to append a 2B CRC to each 32B data chunk. Instead of relying on CRC solely for basic integrity checking, we integrate it into a hybrid RS/CRC scheme, where per-chunk CRCs serve as low-latency filters that detect errors and avoid triggering full RS decoding. This opportunistic use of CRC dramatically reduces read and write amplification, without modifying the HBM physical interface or access granularity.
To organize large-size RS codeword efficiently, we exploit the inherent parallelism of HBM stacks (e.g., each HBM3E stack includes 16 independent channels, each 128 bits wide). We stripe each RS codeword across $s$ such channels to maximize bandwidth and parallel access, as illustrated in Fig. 2. Specifically, each RS codeword contains a number of 32B user data chunks and one or multiple 32B parity chunks. Each 32B chunk, whether data or parity, is appended with its own 2B CRC, forming a 34B unit. These CRC-augmented chunks are striped sequentially across the $s$ channels. Because each chunk is self-contained with its own CRC, this structure supports fine-grained error detection with minimal latency and bandwidth overhead. This hybrid scheme requires no change to the HBM physical interface or access granularity.
@@
Random Write: Fig. 4 illustrates the procedure for handling random write requests. For random writes that update $k$ 32B chunks within a single RS codeword, the memory controller performs a CRC-guided opportunistic partial update to avoid full read-modify-write operations. First, it fetches the current version of the $k$ target chunks along with all the $r$ RS parity chunks. If all CRCs on the fetched chunks pass, the controller proceeds with a partial update as follows. Assume each RS codeword protects a total of $m$ 32B chunks. We construct two $m\times$32B vectors ${\bf D}_{old}$ and ${\bf D}_{new}$: one containing the old values of the $k$ chunks in their original positions and zeros elsewhere, and another containing the new values of those same $k$ chunks (with zeros in the remaining positions). These two sparse vectors are independently RS-encoded to generate two sets of $r$ parity chunks $\text{RS}({\bf D}_{old})$ and $\text{RS}({\bf D}_{new})$. Because RS code is a linear block code, the difference between them captures the net parity update due to the modified data chunks. Accordingly, the new RS parity ${\bf D}_{new}$ is computed by XORing the original parity ${\bf P}_{old}$ with the difference of the two RS-encoded results: $\mathbf{P}_{\text{new}}=\mathbf{P}_{\text{old}}\oplus\text{RS}(\mathbf{D}_{\text{new}})\oplus\text{RS}(\mathbf{D}_{\text{old}})$.
@@
Sequential Read/Write: Sequential read and write requests are handled in a much simpler manner. For sequential reads, we directly fetch the entire RS codeword, including data and parity, and perform RS decoding without first checking CRCs. At high BER, errors are likely, and skipping CRC avoids redundant checks. The RS decoder can naturally early-terminate if no errors are present, minimizing latency and energy. For sequential writes, the controller performs RS encoding and per-chunk CRC encoding in a single pass, then writes the full hybrid RS/CRC codeword directly to HBM.
@@ III-B Importance‑Adaptive ECC Protection
As discussed above, implementing large-codeword RS decoders incurs substantial silicon cost, particularly under multi-TB/s HBM throughput constraints. To mitigate this, we use importance-adaptive ECC that selectively protects only the data bits that matter most for inference quality. Many AI inference workloads, especially with BF16 or FP8, are numerically resilient: mantissa flips often have negligible effect compared to exponent flips. By applying ECC only to the sensitive portions, we cut RS decoder load, ECC traffic, and energy. This mechanism is orthogonal to numeric format: it operates on bit-planes and only requires a per-format criticality map. We report BF16 in experiments for space and prevalence.
To support this functionality efficiently, we adopt a bit-plane–oriented in-memory data placement scheme. Consider a block of $m$ numerical values $\{x_{1},\dots,x_{m}\}$, each $n$ bits wide. Each value $x_{j}$ is represented as a bit vector $b_{j,n-1},b_{j,n-2},\dots,b_{j,0}]$, where $b_{j,i}\in\{0,1\}$ is the $i$-th bit of $x_{j}$. The $i$-th bit-plane is defined as $P_{i}=\{b_{1,i},b_{2,i},\dots,b_{m,i}\}$, i.e., all bits at position $i$ across the block. Let $\mathcal{S}\subseteq\{0,\dots,n-1\}$ denote the set of critical planes that require protection, and let $\gamma=\frac{|\mathcal{S}|}{n}$ denote the protected-plane ratio, with $0\leq\gamma\leq 1$. Only bits in planes $\mathcal{S}$ are processed through CRC and RS encoding; all other bits bypass the ECC logic entirely. Accordingly, we can reduce the RS decoder silicon cost by approximately $1-\gamma$.
@@ IV Evaluation
We use Intel Pin [7] to collect fine-grained LLM inference traces and drive a hybrid evaluation composed of (i) an analytic ECC model (CRC pass probability, RS escalations/read-modify-write amplification, effective bandwidth) and (ii) a modified DRAMSim3 [8] configured as an HBM3-class system with $16\times 128$-bit channels and $\approx$1 TB/s, augmented with hooks for ECC-induced traffic and parameterized encoder/decoder service times.
@@ IV-B Impact of Random Access Intensity
We sweep the random‐access ratio from 0% to 10% at BER $10^{-3}$ (Fig. 6). With purely sequential traffic, throughput rises with codeword size from 13.90 tokens/s (64B) to 18.05 tokens/s (2048B) due to parity amortization. Even small randomness erodes this benefit: at 2% random, 2048B drops to 14.26 tokens/s (-21.0%) while 64B is 13.85 tokens/s (-0.4%). At 10% random, 64B holds 13.64 tokens/s (-1.9%), 256B falls to 11.87 tokens/s (-19.2%), and 2048B collapses to 7.31 tokens/s (-59.5%). The cause is CRC failures on 34B chunks that escalate into full-codeword fetches or read-modify-writes, whose penalty grows with codeword size. For workloads with even modest randomness, moderate codeword sizes (256B–512B) strike the best balance between parity amortization and escalation overhead.